Write the overall cell potential and calculate the standard cell potential of the following

and balance the equation.


Ag⁺ + e → Ag E⁰ = 0.80 V

Mn⁺² + 2e → Mn E⁰ = -1.18 V


1
Expert's answer
2021-01-13T13:57:57-0500

E=E0(Ag)−E0(Mn)=0.80−(−1.18)=1.98V

AnodeMnMn−2e=Mn+2(+)CathodeAgAg++1e=Ag(−)